# Real-time 3D semantic occupancy prediction for autonomous vehicles using memory-efficient sparse convolution

### v1

@@ V-D Real-time Inference
Inference are performed on RTX 4090 with six multi-camera images for the other models at 1600x900 image resolution. Since our model focuses on monocular camera, we use a batch size of 6 for fair comparison. We compare the inference time and GPU memory usage in Table IV, showing a 6 - 10 times increase in inference time, translating to a potential frame rate of 30 FPS and an achievable frame rate of 20 FPS. while utilizing only 1.2GB of GPU memory. For real-time inference demo video, please refer to our supplementary material.
@@ V-E Ablation Study
Referring to Table V, we conducted ablation studies by removing the class balanced cross entropy loss and use standard focal loss, removing the squeeze and excite layer, and removing the higher level feature input tensor. ”Ours” refers to the current model with all three components included. The results are validated against nuScenes validation dataset. The semantic segmentation mIoU and completion IoU are reported. We observed that removing the class balance loss results in a severe drop in both completion IoU and mIoU. This validates the problem of class imbalance in nuScenes dataset, and class balance loss correctly addresses it by leading the model to be sensitive to all classes in both scene completion and semantic segmentation. The incorporation of Squeeze-and-Excite layers leads to slight improvements in both mIoU and completion IoU. This can be attributed to the use of global average pooling, which enhances the model’s adaptability to varying outdoor scenes. Lastly, concatenating higher level image features with the initial $RGB+Intensity$ features achieved no significant performance improvements. This outcome could suggest that the initial feature set was likely sufficient for the task. It could also indicate the feature extractor, EfficientNetV2 [27], might have limited capacity to produce meaningful outdoor scene features.
@@
Description
Semseg mIoU Completion IoU
Ours
0.3603 0.5329
Ours w/o CBLoss
0.3240 0.3911
Ours w/o SELayer
0.3360 0.4983
Ours w/o features
0.3622 0.5110
VI Conclusions and future work
In this paper, we presented a Minkowski engine-based sparse convolution model capable of semantic scene completion on diverse outdoors autonomous vehicle driving scenes using LiDAR and a front-view camera. In particular, we demonstrated that the model is able to perform 3D semantic occupancy prediction with much less computational resources as compared to many current methods, achieving real-time inference, while being on par in accuracy. More broadly, the sparse convolution approach we propose for AVs offers a pathway to enhance the efficiency of real-time robotics perception.
Future work involves several aspects. Firstly, we hope to extend the current model to a multi-view camera setup for 360 degree occupancy prediction of the ego vehicle. Secondly, it is challenging to perform scene completion at distances beyond 20 meters due to the sparsity of LiDAR points (specifically, a 32-beam setup in nuScenes [23]). Incorporating probabilistic techniques to integrate distant camera features into 3D space may help achieve better accuracy. Thirdly, performing self-supervision by creating pseudo dense occupancy ground truth via monocular 2D camera-based depth estimation and semantic segmentation helps to expand the training dataset, allowing the current model to be more robust to complex driving environments. Lastly, on the issue of handling heavily occluded scenes, the integration of a generative adversarial sparse network as a discriminator into the existing scene completion pipeline can be investigated.
APPENDIX
VI-A More Qualitative Examples
VI-B Lambda constant investigation
Using a lambda value of 0.5 yields the best results, with a minor yet notable improvement in Semantic Segmentation mIoU to 0.3555 and a more significant enhancement in Completion IoU to 0.5243.

### v2

@@ V-D Real-time Inference
Inference are performed on RTX 4090 with six multi-camera images for the other models at 1600x900 image resolution. Since our model focuses on monocular camera, we use a batch size of 6 for fair comparison. We compare the inference time and GPU memory usage in Table IV, showing a 6 - 10 times increase in inference time, translating to a potential frame rate of 30 FPS and an achievable frame rate of 20 FPS while utilizing only 1.2GB of GPU memory.
@@ V-E Ablation Study
Referring to Table V, we conducted ablation studies by removing the class balanced cross entropy loss and use standard focal loss, removing the squeeze and excite layer, and removing the higher level feature input tensor. ”Ours” refers to the current model with all three components included. The results are validated against Occ3D-nuScenes validation dataset. The semantic segmentation mIoU and completion IoU are reported. We observed that removing the class balance loss results in a severe drop in both completion IoU and mIoU. This validates the problem of class imbalance in Occ3D-nuScenes dataset, and class balance loss correctly addresses it by leading the model to be sensitive to all classes in both scene completion and semantic segmentation. The incorporation of Squeeze-and-Excite layers leads to slight improvements in both mIoU and completion IoU. This can be attributed to the use of global average pooling, which enhances the model’s adaptability to varying outdoor scenes. Lastly, concatenating higher level image features with the initial $RGB+Intensity$ features achieved no significant performance improvements. This outcome could suggest that the initial feature set was likely sufficient for the task. It could also indicate the feature extractor, EfficientNetV2 [30], might have limited capacity to produce meaningful outdoor scene features.
@@
Description Semseg mIoU Completion IoU
Ours 0.3603 0.5329
Ours w/o CBLoss 0.3240 0.3911
Ours w/o SELayer 0.3360 0.4983
Ours w/o features 0.3622 0.5110
Referring to table VI, using a lambda value of 0.5 yields the best results, with a minor yet notable improvement in Semantic Segmentation mIoU to 0.3555 and a more significant enhancement in Completion IoU to 0.5243.
@@
VI Conclusions and future work
In this paper, we presented a Minkowski engine-based sparse convolution model capable of semantic scene completion on diverse outdoor autonomous vehicle driving scenes using LiDAR and a front-view camera. In particular, we demonstrated that the model is able to perform 3D semantic occupancy prediction with significantly fewer computational resources compared to many current methods, achieving real-time inference while maintaining comparable accuracy. More broadly, the sparse convolution approach we propose for AVs offers a pathway to enhance the efficiency of real-time robotics perception.
To address the challenge of detecting and segmenting small objects, incorporating an object detection preprocessing step on the RGB images can be considered. By utilizing a pretrained lightweight object detection module, the 2D bounding boxes can be projected into the 3D space using the same camera-to-LiDAR transformation matrices, subsequently creating 3D ROI (Region of Interest) masks. These masks complement the input sparse tensor into the current network, providing guidance towards small objects during training and improving their segmentation accuracy.
Future work involves several aspects. Firstly, we aim to extend the current model to a multi-view camera setup for 360-degree occupancy prediction of the ego vehicle. Secondly, performing scene completion at distances beyond 20 meters proves challenging due to the sparsity of LiDAR points, specifically when using a 32-beam setup as in the nuScenes dataset [23]. Incorporating probabilistic techniques to fuse distant camera features and radars into the 3D space may help achieve better accuracy in these regions. Thirdly, employing self-supervision by creating pseudo-dense occupancy ground truth via monocular 2D camera-based depth estimation and semantic segmentation can help expand the training dataset, allowing the current model to become more robust in complex driving environments.